A student reacts 20.0 grams of sodium hydroxide with excess hydrochloric acid, and obtains a yield of 25.0 grams of sodium chloride. What is the percent yield of sodium chloride?

NaOH + HCl --> NaCl + H20
1
Expert's answer
2020-12-06T13:52:54-0500

NaOH + HCl --> NaCl + H20


Molar mass of NaOH = 40 g/mol


Molar mass of NaCl = 58.5 g/mol


As hydrochloric acid is in excess and 20 g of NaOH is present so, NaOH is the limiting reagent .


According to the above equation,

1 mole of NaOH produce 1 mole of NaCl


so,

40 g of NaOH gives 58.5 g of NaCl

1 g of NaOH gives = (58.5/40) g

20 g of NaOH gives = (58.5/40)×20 g

= 29.25 g

Hence, 29.25 gram of NaCl will produce


Theoriticl yield = 29.25 g

Observed yield = 25.0 g


Percentage of yield

= (Observed yield/Theoriticl yield)×100 %

= (25/29.25)× 100 %

= 85.47%


Hence, percentage of yield = 85.47%